A collection of nickels and dimes is worth $3.30. There are 42 coins in all. How many of each kind of coin are there?
this is the equation i wrote: .05x + .10y = 3.30
x + y = 42

Found 2 solutions by josmiceli, stanbon:
Answer by josmiceli(19441)

Correct- now solve for x and y

Answer by stanbon(75887)

A collection of nickels and dimes is worth $3.30. There are 42 coins in all. How many of each kind of coin are there?
this is the equation i wrote:
.05x + .10y = 3.30
x + y = 42
----
Multiply thru 1st Eq. by 100.
Multiply thru 2nd Eq. by 5.
--------------------
5x + 10y = 330 cents
5x + 5y = 5*42
-----
Subtract and solve for "y":
5y = 120
y = 24 (# of dimes)
---
Since x+y = 42, x = 42-24 = 18 (# of nickels)
=================
Cheers,
stan H.